OpenStudy (anonymous):

Shoe sizes and foot length are related by the formula S = 3F - 24, where S represents the shoe size and F represents the length of the foot, in inches. Which of the following is a formula for foot length, F? Answer F equals three S plus twenty-four F equals three S minus twenty-four F equals S plus twenty-four all over three F equals S minus twenty-four all over three

sam (.sam.):

F equals S plus twenty-four all over three

OpenStudy (anonymous):

3F = S+24 F = (1/3)S + 8

OpenStudy (anonymous):

F equals S plus twenty-four all over three